INFRARED PHOTODISSOCIATION SPECTROSCOPY OF FIRST ROW TRANSITION METAL-CARBONYL CATIONS

\underline{ANTONIO D. BRATHWAITE}, ALLEN M. RICKS, ZACH D. REED, MICHAEL A. DUNCAN, Department of Chemistry, University of Georgia, Athens, GA 30602-2256.

Transition metal-carbonyl cations are generated in a laser vaporization/supersonic expansion cluster source, mass selected and studied using infrared laser photodissociation spectroscopy. The carbonyl stretching region (2050-2350 cm^{−1}) is probed using a tunable infrared OPO/OPA system. Several cluster sizes are investigated and insight into their stability and geometry is obtained. Cu(CO)_{4}^{+} has a complete coordination sphere, consistent with 18-electron stability and a tetrahedral structure similar to that of isovalent Ni(CO)_{4}. Ti(CO)_{6}^{+} has a complete coordination sphere and does not satisfy the 18-electron rule. DFT calculations are performed and reported to corroborate the experimental data.